1.     Find out number of endogenous and exogenous variables.

2.     Find out equilibrium level of national income in closed and open economy.

3.     Find out equilibrium level of consumption in closed and open economy.

4.     Find out equilibrium level of savings in closed and open economy.

5.     Prove that S + T = I + G + R in closed economy and S + T + M = I + G + R + X in an open economy.

6.     Find out budget surplus or budget deficit in closed and open economy.

7.     Find out balance of trade or trade deficit or surplus in the economy.

8.     Find out foreign trade multiplier.

9.     Let imports are increased from M to new imports M + (20 + 0.24Y), then find out new equilibrium income in an open economy.

 

 


1
Expert's answer
2021-06-12T10:44:26-0400

National income model be:

"Y = C + I + G"

"C = a + b(Y \u2212 T)" "a > 0, 0 < b < 1"

"T = d + tY" "d > 0, 0 < t < 1"

where Y is national income, C is (planned) consumption expenditure, I is investment expenditure, G is government expenditure and T is taxes

1:) Endogenous: Y, C, T

Exogenous: I, G


2:

0pen economy

"Y = C + I + G + (X \u2013 M)"

closed economy "Y = C + I + G" (Y=national income, C =consumption expenditure, I= investment, G= government expenditure and (X – M)=imports and export's)


3: closed economy

"C = a + b(Y \u2212 T)"

in an open economy

"C = C + bY"

( consumption (C) at a given level of income (Y) is equal to autonomous consumption (C) + b times of given level of income.)

4:"S = I + NX" (closed economy)

"S = Y \u2013 T \u2013 C." (open economy)


5:"Y= C + I + G + (X \u2013 M)"

"Y = C + S + T"

"C + I + G + (X \u2013 M)=C + S + T"

"C-C+I+G+X=S+T+M"

"I+G+X=S+T+M"


6: Budget Deficit = Total Expenditures by the Government − Total Income of the government


7:

Balance of trade "=X-M"


8:"K=\frac{1}{S+M}"


9:"Y = C + I + G + (X \u2013 M)"

"Y = C + I + G + (X \u2013 M+20+0.24Y)"

"Y = C + I + G + X \u2013 M-20-0.24Y"

"Y+0.024Y = C + I + G + X \u2013 M-20"

"1.24Y = C + I + G + X \u2013 M-20"